OpenStudy (anonymous):

Which of the following is the equation of the axis of symmetry of the quadratic equation y = 2x2 – 12x + 22?

OpenStudy (anonymous):

axis x=-b/2a in this case x=-(-12)/2(2)=12/4=3 or x=3